OpenStudy (anonymous):

Identify the vertex, focus, and directrix of the parabola with the equation x^2 – 6x – 8y + 49 = 0

OpenStudy (anonymous):

@thomaster ?

OpenStudy (whpalmer4):

Do you know how to complete the square?

OpenStudy (anonymous):

no..could you please show me?

OpenStudy (whpalmer4):

Okay, \[x^2-6x+n\]is an expression which could be written like \[(x+a)^2\]if we only knew what value to use for \(a\), right? \[(x+a)^2 = (x+a)(x+a) = x^2 + ax + ax + a^2 \]\[\qquad=x^2+2ax+a^2\] If we line those two up over each other: \[x^2-6x+n\]\[x^2+2ax+a^2\] They will be equal if (and only if)\[-6x=2ax\]and\[n=a^2\]

OpenStudy (anonymous):

umm..ok, them what do we need to do?

OpenStudy (whpalmer4):

so, our job is to figure out how much to take from that 49 we have sitting there in the equation so that we can write our equation as something like \[(x+a)^2 = 8y+b\] So, \[-6x=2ax\]what does \(a=\)

OpenStudy (anonymous):

umm 2ax? im not sure..how do you find a?

OpenStudy (whpalmer4):

\[-6x=2ax\]Solve for \(a\) Do you see any common factors you can cancel?

OpenStudy (anonymous):

x?

OpenStudy (anonymous):

ok, so -6=2a?

OpenStudy (whpalmer4):

Yes, good so far, what is the value of \(a\) if that is the case?

OpenStudy (anonymous):

im not sure, how would you find it? we know that 2a=-6...?

OpenStudy (anonymous):

i just dont know what to do next..

OpenStudy (whpalmer4):

Uh, to solve for \(a\) in that equation, wouldn't you divide both sides by the number that appears in front of the \(a\)?

OpenStudy (whpalmer4):

What number times 2 gives you -6? That number is \(a\).

OpenStudy (anonymous):

ooohh yeah! ok, so a would equal 3

OpenStudy (anonymous):

wow sorry blond moment:P

OpenStudy (whpalmer4):

okay, so we have \(a = -3\) what is \(a^2=\)

OpenStudy (anonymous):

oh i forgot about the negative... so it would be -3

OpenStudy (anonymous):

a^2=9

OpenStudy (anonymous):

then what?

OpenStudy (whpalmer4):

Ok, so we've found that we can take \(x^2-6x\) and add \(9\) to it to get a "perfect square" which we can write as \((x-3)^2\): \[(x-3)^2=(x-3)(x-3) = x(x-3)-3(x-3) = x*x-3x-3x-3(-3)\]\[\qquad = x^2-6x+9\] Any question about that?

OpenStudy (whpalmer4):

(other than "why would you want do to THAT?!?" :-)

OpenStudy (anonymous):

haha:) umm so you just expand it in order to simplify it?

OpenStudy (whpalmer4):

We actually want to keep in the compact form we have: \((x+3)^2\) But let's take a minute and see just what we did: We had \[x^2-6x+\text{<stuff not involving x>}\] We found that if we took half of the number in front of \(x\) (call it \(a\)), and squared it (\(a^2\)), we could write \((x+a)^2\) instead of \(x^2-6x\) However, we need to make an adjustment. We're taking \(a^2=9\) away from the rest of the equation, so we can't just write\[(x-3)^2-8y+49=0\]and expect good things to happen. Instead, we need to balance the equation so that it is still a valid equation. To write \((x-3)^2\) we needed another \(a^2=9\). If we just add that value to both sides of the equation, then we're in the clear, because you can add the same quantity to both sides of an equation without disturbing the balance. \[x^2-6x -8y + 49 = 0\]\[x^2-6x +9 -8y + 49 = 0+9\]\[(x^2-6x+9)-8y+49=9\]\[(x-3)^2-8y+49=9\]Do those steps make sense to you?

OpenStudy (whpalmer4):

Think of it as making neat square arrays of blocks. You've got some blocks already on the table, but you need a few more to complete your square. The teacher wants to keep things fair, so that you don't get more blocks than the kid next to you, so when you raise your hand and say "I need 9 more blocks to complete my square", the teacher comes over and gives you 9 more blocks, and gives the kid next to you 9 more blocks. You've still both got the same number of blocks as each other, but now you can arrange yours in a square.

OpenStudy (anonymous):

oh, ok:) well what do we do next?

OpenStudy (anonymous):

or are we done?

OpenStudy (whpalmer4):

So, we've completed the square. Our work is just beginning :-)

OpenStudy (anonymous):

ooohh..yey..:)

OpenStudy (whpalmer4):

what forms of the parabola equation have you learned so far?

OpenStudy (whpalmer4):

have you heard of vertex form?

OpenStudy (anonymous):

yes, that and standard form..

OpenStudy (whpalmer4):

Okay, does what we have here: \[(x-3)^2 -8y + 49=9\]look like it would be easy to put into vertex form?

OpenStudy (anonymous):

well vertex form is y=a(x-h)^2+k...and we dont know what (x-h)^2 would be..

OpenStudy (whpalmer4):

Why don't you try solving what we have for \(y\)

OpenStudy (anonymous):

its -8y

OpenStudy (whpalmer4):

Take the equation we have, and use algebra to rearrange it so we have \[y=\text{<stuff>}\]

OpenStudy (whpalmer4):

\(y\) will not appear in <stuff> the only \(y\) will be all alone on the left side

OpenStudy (whpalmer4):

\[(x-3)^2-8y+49=9\]Add \(8y\) to both sides\[(x-3)^2-8y+49+8y=9+8y\]\[(x-3)^2+49=9+8y\]Any question about that?

OpenStudy (whpalmer4):

Next, flop the equation: \[9+8y=(x-3)^2+49\] Any question about that?

OpenStudy (anonymous):

no questions:) whats next?:)

OpenStudy (whpalmer4):

What's next is I would like you to complete the process of isolating \(y\)

OpenStudy (whpalmer4):

You're going to have to subtract 9 from both sides, and then divide both sides by some number

OpenStudy (anonymous):

ok, so we would have 8y-(x-3)^2+40...and now we divide by 8 on each side right?

OpenStudy (whpalmer4):

what happened to the = sign?

OpenStudy (anonymous):

oh, whoops..8y=(x-3)^2+40

OpenStudy (whpalmer4):

that's better. Now divide both sides by 8

OpenStudy (anonymous):

ok...y=(x-3)^2+5

OpenStudy (whpalmer4):

yes, does that look like vertex form?

OpenStudy (anonymous):

yes:)

OpenStudy (whpalmer4):

okay, so what is the vertex?

OpenStudy (anonymous):

(x-3)^2

OpenStudy (anonymous):

right?

OpenStudy (whpalmer4):

no. the vertex is a point

OpenStudy (whpalmer4):

vertex form: \[y = a(x-h)^2+k\]vertex is at \((h,k)\)

OpenStudy (whpalmer4):

No!

OpenStudy (whpalmer4):

Compare equations: \[y = a(x-h)^2+k\]\[y=(x-3)^2+5\]

OpenStudy (whpalmer4):

what is the value of \(k\)?

OpenStudy (anonymous):

im not sure?

OpenStudy (anonymous):

oh! its 5..

OpenStudy (whpalmer4):

Yes, \(k=5\). What is the value of \(h\)?

OpenStudy (whpalmer4):

Well, it appears you left sometime during the time that OpenStudy was taking its usual mid-morning nap. The vertex is at \((h,k)\) once you figure out the value of \(h\). Note very carefully that the formula is \[y = a(x-h)^2+k\]so make sure you get the proper sign!

OpenStudy (whpalmer4):

We have a parabola with \(x^2\) in it, so it opens either up or down. A parabola with \(y^2\) in it would open either right or left. If the parabola opens up or down, the directrix will be a line with a constant \(y\) value. The directrix and the focus are located equidistant from the vertex.